Use the following sales data to answer the questions. Month                              Sales January&nbsp

  1. Use the following sales data to answer the questions.

Month                              Sales

January                              $250,000

February                            $200,000

March                               $300,000

April                                  $350,000

May                                   $450,000

  1. Using a two month moving average, what are the expected sales for June?
  1. Using a three month moving average, what are the expected sales for June?
  1. Ed Rogers owns an appliance store. Sales data on a particular model of a DVD player for the past six months are shown below, along with the results of two different forecasting models:

Month

Sales

Forecast 1

Forecast 2

Jan

35

30

33

Feb

29

28

32

Mar

39

43

35

Apr

42

40

45

May

51

48

52

Jun

56

55

52

Which is the better forecasting model, based on the MAD criterion?

  1. Ed Rogers owns an appliance store. Sales data on a particular model of a DVD player for the past six months are:

Month

Sales

Jan

35

Feb

29

Mar

39

Apr

42

May

51

Jun

56

Forecast sales for July using an exponential smoothing model with a smoothing constant of 0.40. Assume that the forecast for May was 36.25.
Answer #1

Qs 1)

a) A two period moving average method averages the actual value for the previous two periods to generate the forecast for the next period. This can be calculated as the sum of the actual value for the previous two periods/2

So using the above formula, expected sales for June = (350000+450000) / 2 = 800000/2 = 400000

b) A three period moving average method averages the actual value for the previous three periods to generate the forecast for the next period. This can be calculated as the sum of the actual value for the previous three periods/3

So using the above formula, expected sales for June = (300000+350000+450000) /3 = 1100000/3 = 366666.67